x^2+6x+9=25, give exact and approximate within 3 decimal places.
Answer by ewatrrr(24785)

 
Hi
x^2+6x+9=25
factoring left side
(x+3)^2 = 25
x+3 = ± 5 |taking the square root of both sides of the EQ
x = -3 ± 5
x = -8 and x = 2